# A Scalable Quantum Non-local Neural Network for Image Classification

### v1

@@ Abstract
Non-local operations play a crucial role in computer vision enabling the capture of long-range dependencies through weighted sums of features across the input, surpassing the constraints of traditional convolution operations that focus solely on local neighborhoods. Non-local operations typically require computing pairwise relationships between all elements in a set, leading to quadratic complexity in terms of time and memory. Due to the high computational and memory demands, scaling non-local neural networks to large-scale problems can be challenging. This article introduces a hybrid quantum-classical scalable non-local neural network, referred to as Quantum Non-Local Neural Network (QNL-Net), to enhance pattern recognition. The proposed QNL-Net relies on inherent quantum parallelism to allow the simultaneous processing of a large number of input features enabling more efficient computations in quantum-enhanced feature space and involving pairwise relationships through quantum entanglement. We benchmark our proposed QNL-Net with other quantum counterparts to binary classification with datasets MNIST and CIFAR-10. The simulation findings showcase our QNL-Net achieves cutting-edge accuracy levels in binary image classification among quantum classifiers while utilizing fewer qubits.
Codehttps://anonymous.4open.science/r/QNL-Net/
@@ Introduction
Computer vision has become a cornerstone of artificial intelligence, consisting of a wide array of applications such as autonomous driving (Bao et al. 2023), medical imaging (Li et al. 2023), healthcare (Zhou et al. 2022), etc. An essential task in this domain is image classification, where the goal is to assign a label to an image based on its visual significance. This task also builds upon more complex applications such as image segmentation (Minaee et al. 2021), object detection (Amjoud and Amrouch 2023; Fu et al. 2023), and scene understanding (Naseer, Khan, and Porikli 2018). Image classification approaches have been significantly advanced by Convolutional Neural Networks (CNNs) (O’shea and Nash 2015), which achieve state-of-the-art performance on datasets like ImageNet (Krizhevsky, Sutskever, and Hinton 2012). However, CNNs are limited by their local receptive fields, which restricts them from capturing broader contextual information and long-range dependencies within an image. To overcome these limitations, non-local neural networks were introduced to capture long-range dependencies in data, an extension of the self-attention mechanism used in Transformer architectures (Vaswani et al. 2017). Initially proposed for computer vision applications by (Wang et al. 2018), non-local neural networks capture global context effectively and have shown significant improvements in computer vision tasks that benefit from modeling long-range dependencies.
@@
Parallel to these advancements, quantum machine learning (QML) (Biamonte et al. 2017) has emerged as a revolutionary technology built upon the principles of quantum mechanics, which describes the behavior and nature of atoms at the smallest fundamental level and applies them to machine learning. In classical computing, information only exists in bits, which are 0 or 1. In contrast, quantum computing introduces the concept of qubits that can have quantum states |$\ket{0}$ and |$\ket{1}$ simultaneously, taking advantage of the concept of superposition and introducing features such as quantum parallelism (Nielsen and Chuang 2001). Theoretically, this allows for speeding up computations and devising algorithms that can solve complex challenges more efficiently compared to classical computing where computation might be expensive and inefficient, and potentially revolutionize fields like pattern recognition and image classification, optimization (Abbas et al. 2023), cryptography (Bernstein and Lange 2017), etc. QML also introduces algorithms for classification problems such as QSVM, Quantum Kernel methods, Variational Quantum Classifiers (VQC), etc. VQC is a particularly interesting approach within QML because it lets us combine classical and quantum computing to use a quantum circuit as the core algorithm consisting of quantum gates that can be parameterized (Benedetti et al. 2019; Peddireddy, Bansal, and Aggarwal 2023). These parameters can be optimized using classical methods, enabling the model to be trained with a hybrid quantum-classical approach, which we utilize in our work.
@@
Researchers have still been trying to understand whether QML offers a significant advantage to classical machine learning theoretically and practically, and it has been an active field of research for long (Biamonte et al. 2017). A sub-domain in this field, Quantum Neural Networks (QNNs), has been developed intensively to determine whether these are capable of outperforming classical neural networks. It was also recently explored in (Abbas et al. 2021) by performing simulations on actual quantum hardware and proving that QML does, in fact, offer several advantages. However, a few challenges exist due to insufficient capabilities for fault-tolerance, still-evolving quantum error correction techniques, and quantum scalability, which presents an avenue for revolutionary research in this field and transitioning to a new era of quantum computing from the present noisy intermediate-scale quantum (NISQ) era (Preskill 2018; Gujju, Matsuo, and Raymond 2024). The present challenges limit the practical applications of QML algorithms. Still, despite that, developing QML research and devising new quantum algorithms offers an outlet to test methodologies as a proof-of-concept on a smaller scale and deploy applications in the future while we approach scalable quantum processors.
In this context, our paper proposes a Quantum Non-local Neural Network (QNL-Net), which combines the principles of quantum computing with the non-local neural network mechanism. By leveraging quantum entanglement, the QNL-Net can establish non-local correlations between qubits, mimicking the behavior of classical non-local operations while exploiting the advantages of quantum mechanics to enhance machine learning performance and capabilities. This paper aims to improve pattern recognition and binary classification tasks in computer vision by capturing long-range dependencies more effectively. The main contributions of this paper are:
@@
    We introduce a scalable and customizable quantum implementation of the non-local neural network mechanism that utilizes only four qubits to achieve state-of-the-art performance on binary classification tasks in computer vision.
@@
    We also utilize classical machine learning techniques for dimensionality reduction of features before these are processed by the QNL-Net framework, therefore leveraging the advantages of hybrid quantum-classical machine learning models.
@@
    We comprehensively evaluate the QNL-Net’s performance on benchmark datasets such as MNIST and CIFAR-10, showcasing its robustness to noise and its potential advantages over traditional quantum binary classification models. Our hybrid classical-quantum models, particularly the CNN-QNL-Net, achieved near-perfect binary classification accuracies on the MNIST dataset (99.96% test accuracy) and outperformed benchmark quantum classifiers such as QTN-VQC (98.6% with 12 qubits) (Qi, Yang, and Chen 2023) and Hybrid TTN-MERA (99.87% with 8 qubits) (Grant et al. 2018) using significantly fewer qubits, i.e., 4 qubits.
@@
Image classification using quantum machine learning (QML) became an area of significant interest due to its potential advantages over classical methods. One notable approach was Quantum Convolutional Neural Networks (QCNNs), which utilized quantum circuits to implement convolutional operations, focusing on quantum phase recognition and quantum error correction optimization techniques (Cong, Choi, and Lukin 2019). Another significant development was the Quanvolutional Neural Network, introduced by (Henderson et al. 2020) in 2019, which employed quantum convolution (quanvolutional) layers. These layers transformed classical data using random quantum circuits to extract features, akin to the feature extraction process in classical CNNs. Their results demonstrated superior accuracy and training performance compared to classical CNNs. QML models have also shown efficacy in binary classification tasks for noisy datasets and images (Schetakis et al. 2022). Additionally, recent work by (Cherrat et al. 2024) developed an approach for loading matrices as quantum states and introducing trainable quantum orthogonal layers adaptable to different quantum computer capabilities. This method yielded promising results on superconducting quantum computers.
Several works also explored the application of quantum neural networks for binary classification tasks in computer vision, providing benchmarks for our results. QTN-VQC (Qi, Yang, and Chen 2023) built a framework with quantum circuits for tensor-train networks integrated with variational quantum circuits for an efficient training pipeline. Another work introduced hierarchical quantum classifiers (Grant et al. 2018), which utilized several expressive circuits to classify highly entangled quantum states and demonstrated robustness to noise. A scalable approach for quantum neural networks (SQNNs) for classification was discussed in (Wu, Tao, and Li 2022), where authors proposed a strategy to use multiple small-scale quantum devices to extract local features and perform prediction over these collected features. (Jiang, Xiong, and Shi 2021) presented the QuantumFlow model, which represented data as unitary matrices to achieve quantum advantage, reducing the cost complexity of unitary matrices-based neural computation. These recent advancements in QML models demonstrated robustness in image classification, making them suitable for handling higher-dimensional data more effectively than their classical counterparts.
@@
Traditional convolution operations in convolutional neural networks (CNNs), a popular choice for computer vision models, process a local neighborhood, and applying these operations repeatedly to capture long-range dependencies causes incremental growth in the receptive field. This has several drawbacks associated with computational inefficiency and difficulties in optimization. Non-local neural networks address these limitations by introducing non-local operations that compute the response at a position as a weighted sum of the features at all positions in the input. These are simple operations that are highly efficient and generic in capturing long-range dependencies, which is of utmost importance in computer vision (Wang et al. 2018).
A generic non-local operation for an input signal’s (image, sequence, video) feature map $\mathrm{x}\in\mathbb{R}^{N\times C}$, where $N$ is the number of positions (i.e. pixels) and $C$ is the number of channels, can be defined as:
@@
    where $\theta(\mathrm{x}_{i})=W_{\theta}\mathrm{x}_{i}$ and $\phi(\mathrm{x}_{j})=W_{\phi}\mathrm{x}_{j}$ are linear embeddings. This also relates to self-attention (as explored in (Vaswani et al. 2017)), which non-local networks are just an extension of in the computer vision domain, or more specifically, a generic space or spacetime domain. This is due to the fact that for any $i$, $\frac{1}{\mathcal{C}(\mathrm{x})}f(\mathrm{x}_{i},\mathrm{x}_{j})$ becomes the softmax computation along the dimension $j$.
@@
    $f(\mathrm{x}_{i},\mathrm{x}_{j})=ReLU(w_{f}^{T}[\theta(\mathrm{x}_{i})^{T},\phi(\mathrm{x}_{j})]),\hskip 2.84526pt\mathcal{C}(\mathrm{x})=N;$
A non-local block in spacetime encapsulates the non-local operation in eq.(1) elegantly and can be defined as:
@@
In this work, we introduce the Quantum Non-Local Neural Network (QNL-Net), which utilizes trainable quantum circuits to implement non-local operations, effectively capturing and processing long-range dependencies in input data. The QNL-Net module integrates with classical dimensionality reduction techniques to function as a hybrid quantum-classical classifier. In this section, we first delve into the design and implementation of the QNL-Net module. Next, we discuss the integration of classical dimensionality reduction techniques with the QNL-Net to create a hybrid classifier, highlighting the CNN-QNL-Net and PCA-QNL-Net models. Finally, we cover the post-QNL-Net classical computation.
@@
The QNL-Net mechanism translates classical non-local operations into quantum circuits, enabling the network to exploit the parallelism and entanglement properties of quantum computing. This translation involves designing quantum gates and circuits that can replicate the functionality of classical non-local layers, allowing the network to analyze complex data structures more efficiently.
@@
Figure 1: The Quantum Non-local Neural Network (QNL-Net) mechanism’s four-qubit circuit composed of three parts: (i) Encoder: uses Qiskit’s ZFeatureMap (Kanazawa et al. 2023) to encode classical data into quantum states. (ii) Variational Quantum Circuit (VQC): the classically trainable quantum circuit. (iii) Measurement: the circuit is measured at qubit 0 in the Pauli-$Z$ basis. The encoder and the VQC ansatz have a depth of $r$ and $D$ respectively.
To encode the classical data $\mathcal{X}\in\mathbb{R}^{d}$ into the quantum space where $\mathcal{X}=[y_{0},y_{1},\cdots,y_{k}]$, we utilize Qiskit’s ZFeatureMap (Kanazawa et al. 2023) (as one can see in Fig. 1) which takes advantage of the quantum-enhanced feature space, providing a quantum advantage to classification problems (Havlíček et al. 2019). The feature map $\Phi$ acts as $\Phi:\mathcal{X}\rightarrow\ket{\psi_{\Phi}}$, where the feature space |$\ket{\psi_{\Phi}}$ is an $n$-qubit Hilbert space, such that
|$\ket{\psi_{\Phi}}=(U_{\Phi}(y_{k})H^{\otimes n})^{r}\ket{0^{n}},\hskip 8.53581pt0\leq k\leq N-1,$ (3)
where $r$ is the depth or the number of repetitions of the feature map, $H^{\otimes n}$ is a layer of Hadamard gates acting on all $n$ qubits, and $N$ denotes the number of input vectors in a data sample. $U_{\Phi}$ is the encoding ansatz used in the feature map composed of single-qubit phase-shift ($P_{\lambda}$) gates for rotation about the $Z$-axis,
$U_{\Phi}(\mathcal{X})=\bigotimes^{n}_{k=1}{P_{2*y_{k}}}.$ (4)
@@
Figure 2: The three ansatzes used as the variational quantum circuits (VQCs) in the QNL-Net mechanism. $R_{x},R_{y},R_{z}$ rotation gates represent single-qubit rotations along the $x,y,z$-axes respectively with trainable parameters. The specific strategies for performing entanglement using CX gates are: cyclic pattern (Ansatz-0), reverse linear chain (Ansatz-1), and a mixed pattern (Ansatz-2).
@@
We utilize variational quantum circuits in this mechanism, which can be trained classically to optimize trainable parameters using a predefined cost function (Benedetti et al. 2019). This hybrid quantum-classical approach leverages the expressive power of quantum circuits while making use of classical optimization techniques for efficient parameter tuning. We implement three different ansatzes with five parameterized rotation gates in each, enabling fine-grained control over the quantum state. These three ansatz each have three $CX$ gates with distinct configurations, introducing entanglement-induced correlations between the qubits and ensuring that the quantum state captures the non-local dependencies in the data. The quantum circuits for the ansatzes are presented in Fig. 2.
We initially apply rotation gates to each qubit in all the ansatzes to obtain the following quantum state.
|$\ket{\psi_{1}}=R_{z}(x_{0})_{q_{0}}R_{y}(\theta_{0})_{q_{1}}R_{y}(\phi_{0})_{q_{2}}R_{x}(g_{0})_{q_{3}}\ket{\psi_{\Phi}}$ (5)
@@
The variation in the ansatzes comes from different entanglement strategies described below. The configurations ensure that all qubits are entangled with each other in order to capture the non-local dependencies throughout the circuit. In Ansatz-0, the CX gates form a cyclic pattern, creating a loop of entanglement among the qubits. Ansatz-1’s CNOT gates form a reverse linear chain, creating a backward sequential entanglement. Ansatz-2 forms a non-linear and unique mixed pattern entanglement strategy using the CNOT gates. The resultant quantum states for the respective ansatz are as follows.
|$\ket{\psi_{2}^{[0]}}=CX_{q1q2}CX_{q2q3}CX_{q3q0}\ket{\psi_{1}},$ (6)
|$\ket{\psi_{2}^{[1]}}=CX_{q3q2}CX_{q2q1}CX_{q1q0}\ket{\psi_{1}},$ (7)
|$\ket{\psi_{2}^{[2]}}=CX_{q1q3}CX_{q3q2}CX_{q2q0}\ket{\psi_{1}}.$ (8)
@@
This sequence of gates and entanglements in the ansatzes constitutes one layer of the respective ansatz and can also be represented by the unitary operator, such that,
@@
Now, each layer can be applied for a depth of $D$ to enhance the expressiveness of the model, and therefore, we obtain the final state of the circuit,
|$\ket{\psi_{final}}=(U_{a}(\theta))^{D}\ket{\psi_{\Phi}},$ (11)
where $a$ is the desired ansatz $a$ = 0, 1, or 2. According to the Born rule, measuring any quantum state in the Pauli-$Z$ basis ($\sigma_{z}$) either collapses into the state |$\ket{0}$ with probability $|\alpha|^{2}$ or into the state |$\ket{1}$ with probability $|\beta|^{2}$ (Nielsen and Chuang 2001). In general, the expectation of any observable $\hat{O}$ for a state |$\ket{\psi}$ can be denoted as
@@
$O=\sum^{K}_{k=1}\alpha_{k}P_{k},\hskip 8.53581pt\alpha_{k}\in\mathbb{R},$ (13)
@@
where $P_{k}\in\{I,X,Y,Z\}^{\otimes n}$ denotes a Pauli observable. In our mechanism, we measure the circuit in the Pauli-$Z$ computational basis at only one qubit, $q_{0}$, whereas the rest of the qubits are measured using an Identity ($I$) operation. So, we can denote the measurement of $q_{0}$ for an input state of |$\ket{\psi_{\Phi}}$ as $\braket{Z_{x}}_{\Phi}$, where $x$ is the data embedding on qubit $q_{0}$, and
$\braket{Z_{x}}_{\Phi}=\bra{\psi_{\Phi}}U_{x}^{\dagger}(\theta_{x})Z_{1}U_{x}(\theta_{x})\ket{\psi_{\Phi}},$ (14)
where $\theta_{x}$ is a parameter on the quantum ansatz $U_{x}$ and $Z_{1}$ represents Pauli-$Z$ measurement on the first qubit.

### v2

@@ Abstract
Non-local operations play a crucial role in computer vision, enabling the capture of long-range dependencies through weighted sums of features across the input, surpassing the constraints of traditional convolution operations that focus solely on local neighborhoods. Non-local operations typically require computing pairwise relationships between all elements in a feature set, leading to quadratic complexity in terms of time and memory. Due to the high computational and memory demands, scaling non-local neural networks to large-scale problems can be challenging. This article introduces a hybrid quantum-classical scalable non-local neural network, the Quantum Non-Local Neural Network (QNL-Net), to enhance pattern recognition. The proposed QNL-Net relies on inherent quantum parallelism to allow the simultaneous processing of a large number of input features, enabling more efficient computations in quantum-enhanced feature space and involving pairwise relationships through quantum entanglement. We benchmark our proposed QNL-Net with other quantum counterparts for binary classification with the MNIST and CIFAR-10 data sets. The simulation findings showcase our QNL-Net, which achieves cutting-edge accuracy levels in binary image classification among quantum classifiers while utilizing fewer qubits.
Codehttps://github.com/sparshgup/QNL-Net
@@ Introduction
Computer vision, a fundamental component of Artificial Intelligence (AI), encompasses various applications such as autonomous driving (Bao et al. 2023), medical imaging (Konar et al. 2024), and healthcare (Zhou et al. 2022). Image classification is a crucial task within computer vision, with the aim of assigning labels to images based on their visual content. This task serves as the foundation for more complex applications like image segmentation, object detection, and scene understanding. Convolutional Neural Networks (CNNs) have significantly advanced image classification, achieving top-tier performance on datasets like ImageNet (Krizhevsky, Sutskever, and Hinton 2012). However, CNNs are restricted by their local receptive fields, limiting their ability to capture extensive contextual information and long-range dependencies in images. To address these limitations, non-local neural networks were introduced to capture long-range dependencies, building upon the self-attention mechanism seen in Transformer architectures (Vaswani et al. 2017). Non-local neural networks have proven effective in capturing global context and improving performance in computer vision tasks that require modeling long-range dependencies (Wang et al. 2018). These networks have shown significant improvements in tasks such as image classification, where understanding broader contextual information is crucial for accurate labeling (Shi, Zhao, and Wang 2021). Moreover, non-local operations typically require computing pairwise relationships between all elements in a feature set, leading to quadratic complexity in terms of time and memory. Due to the high computational and memory demands, scaling non-local neural networks to large-scale problems is a challenging proposition.
@@
Quantum Machine Learning (QML) has emerged as a revolutionary technology that combines principles of quantum mechanics with Machine Learning (ML), offering transformative potential in various fields (Biamonte et al. 2017). Unlike classical computing that relies on bits representing |$\ket{0}$ and |$\ket{1}$, Quantum Computing (QC) introduces qubits that can exist in quantum states simultaneously, leveraging superposition and quantum parallelism for enhanced computational capabilities (Nielsen and Chuang 2001). This quantum advantage allows the development of algorithms that can efficiently tackle complex challenges, potentially revolutionizing areas such as pattern recognition, image classification, optimization, and cryptography (Bernstein and Lange 2017). Within QML, algorithms such as Quantum Support Vector Machine (QSVM), Quantum Kernel methods, and Variational Quantum Classifiers (VQC) have been introduced for classification tasks (Benedetti et al. 2019). VQC, in particular, stands out as it combines classical and quantum computing utilizing a quantum circuit with parameterized quantum gates. These parameters can be optimized using classical methods, enabling a hybrid quantum-classical approach to effectively train models (Ding 2024).
@@
Researchers have actively investigated the potential advantages of QML over classical machine learning, both theoretically and practically (Biamonte et al. 2017). A specific sub-domain within this field, Quantum Neural Networks (QNN), has been extensively developed to evaluate their ability to surpass classical neural networks. Recent studies have used simulations on actual quantum hardware to demonstrate the benefits offered by QML (Abbas et al. 2021). Despite these advances, challenges persist in QML, including limitations in fault-tolerance, evolving quantum error correction techniques, and issues related to quantum scalability (Preskill 2018; Gujju, Matsuo, and Raymond 2024). These challenges present opportunities for innovative research in transitioning from the current Noisy Intermediate-Scale Quantum (NISQ) era to a new era of quantum computing (Huang et al. 2021). The limitations in fault-tolerance and quantum scalability impede the practical applications of QML algorithms. However, ongoing research efforts to develop new quantum algorithms and address these challenges provide a platform to test methodologies on a smaller scale and pave the way for future applications as quantum processors become more scalable.
In this study, we introduce a novel Quantum Non-local Neural Network (QNL-Net) that merges QC principles with the non-local neural network mechanism to tackle pairwise non-local operations in a feature set harnessing the inherent characteristics of quantum entanglement. The proposed QNL-Net establishes non-local correlations through quantum-enhanced features, emulating classical non-local operations while leveraging quantum entanglement to enhance model performance and capabilities. The proposed QNL-Net architecture introduces a novel approach that efficiently utilizes fewer qubits, a crucial aspect in the NISQ era. Beyond this efficiency, the architecture’s significance lies in its treatment of rotations around axes and entanglement. The selection of rotation gates in the proposed QNL-Net ansatzes is closely linked to the core concept of non-local neural networks, which aim to capture intricate spatial dependencies within data. In the quantum realm, rotation gates play a pivotal role in achieving this objective by maneuvering quantum states around various axes, thereby facilitating spatial transformations.
The primary objective of this research is to improve the recognition of patterns and binary classification tasks in computer vision by effectively capturing long-range dependencies. QML holds the potential to overcome several drawbacks of classical non-local neural nets, leveraging the unique characteristics of QC. Quantum parallelism allows for the simultaneous processing of a large number of states, enabling more efficient computations in tasks involving pairwise relationships. Moreover, quantum feature mapping in the proposed QNL-Net transforms classical data into high-dimensional quantum spaces where patterns become more interpretable. This helps in understanding the interactions captured by non-local operations. The main contributions of our work are as follows:
@@
    We introduce a scalable Quantum Non-local Neural Network (QNL-Net) to tackle pairwise non-local operations in a feature set harnessing the inherent characteristics of quantum entanglement.
@@
    We also employed classical machine learning techniques for dimensionality reduction of features before these are processed by our proposed QNL-Net framework, therefore leveraging the advantages of hybrid quantum-classical machine learning models.
@@
    Our hybrid classical-quantum models, especially CNN-QNL-Net, achieved high nineties classification accuracy on the MNIST dataset (99.96% test accuracy) and outperformed the benchmark quantum classifiers QTN-VQC ($98.6\%$ with $12$ qubits) (Qi, Yang, and Chen 2023) and hybrid TTN-MERA ($99.87\%$ with $8$ qubits) (Grant et al. 2018) using significantly fewer qubits ($4$ qubits).
@@
Image classification using Quantum Machine Learning (QML) has attracted significant attention due to its potential advantages over classical methods. One notable approach is Quantum Convolutional Neural Networks (QCNN), which utilize quantum circuits for convolutional operations, emphasizing quantum phase recognition and error correction optimization techniques (Cong, Choi, and Lukin 2019). Another significant development is the Quanvolutional Neural Network, which introduces quantum convolution layers to transform classical data using random quantum circuits for feature extraction, which showcases superior accuracy and training performance compared to classical CNNs (Henderson et al. 2020). QML models have shown effectiveness in binary classification tasks for noisy datasets and images (Schetakis et al. 2022). Nevertheless, recent advances include notable work that loads matrices as quantum states and introduces trainable quantum orthogonal layers adaptable to various quantum computer capabilities, showing promising results on superconducting quantum computers (Cherrat et al. 2024). These developments highlight the potential of QML to improve image classification systems.
Various studies have explored the application of Quantum Neural Networks (QNN) in binary classification tasks within computer vision, providing benchmarks for research outcomes. For example, QTN-VQC (Qi, Yang, and Chen 2023) integrated quantum circuits for tensor-train networks with variational quantum circuits to establish an efficient training pipeline. Furthermore, hierarchical quantum classifiers use expressive circuits to classify highly entangled quantum states and demonstrate resilience to noise (Grant et al. 2018). Konar et al. (2023) proposes a scalable approach for Spiking Quantum Neural Networks (SQNN) for classification, suggesting the utilization of multiple small-scale quantum devices to extract local features and make predictions. (Jiang, Xiong, and Shi 2021) presented a Quantum Flow model that represented data as unitary matrices to achieve a quantum advantage, reducing the cost complexity of unitary matrices-based neural computation. These recent advances in QML models have shown robustness in image classification, which is proven effective in handling higher-dimensional data more efficiently than its classical counterparts.
@@
Classical convolution operations in Convolutional Neural Networks (CNN) are widely used in computer vision models but face challenges related to processing long-range dependencies due to the incremental growth in the receptive field with repeated operations. This leads to computational inefficiency and optimization difficulties. To address these limitations, classical Non-local Neural Networks (NL-Net) have been introduced, which incorporate non-local operations that compute responses by aggregating features from all positions in the input, enabling efficient capture of long-range dependencies crucial in computer vision tasks (Wang et al. 2018).
A generic non-local operation for an input signal’s (image, sequence, video) feature map $\mathrm{x}\in\mathbb{R}^{N\times C}$, where $N$ is the number of positions (i.e., pixels) and $C$ is the number of channels, is defined as:
@@
    where $\theta(\mathrm{x}_{i})=W_{\theta}\mathrm{x}_{i}$ and $\phi(\mathrm{x}_{j})=W_{\phi}\mathrm{x}_{j}$ are linear embeddings. This also relates to self-attention (Vaswani et al. 2017), in which NL-Net is just an extension of the computer vision domain, or more specifically, a generic space or spacetime domain. This is due to the fact that for any $i$, $\frac{1}{\mathcal{C}(\mathrm{x})}f(\mathrm{x}_{i},\mathrm{x}_{j})$ becomes the softmax computation along the dimension $j$.
@@
    $f(\mathrm{x}_{i},\mathrm{x}_{j})=ReLU(w_{f}^{T}[\theta(\mathrm{x}_{i})^{T},\phi(\mathrm{x}_{j})]),\mathcal{C}(\mathrm{x})=N;$
A non-local block in spacetime encapsulates the non-local operation in Eq.  (1) elegantly and is defined as:
@@
In this study, we present the Quantum Non-Local Neural Network (QNL-Net), which leverages trainable quantum circuits to implement non-local operations, effectively capturing and processing long-range dependencies within input data. Our proposed QNL-Net architecture integrates classical dimensionality reduction techniques to operate as a hybrid quantum-classical classifier. This section entails the design and implementation of the QNL-Net architecture, followed by the incorporation of classical dimensionality reduction methods to create hybrid classifiers, specifically highlighting the CNN-QNL-Net and PCA-QNL-Net models. Additionally, we explore post-QNL-Net classical computation to enhance the classification process further.
@@
The proposed QNL-Net introduces a novel approach by translating classical non-local operations into quantum circuits, harnessing the parallelism and entanglement properties inherent in quantum computing. This translation process involves the design of quantum gates and circuits that mimic the functionality of classical non-local layers, enabling the network to analyze complex data structures more efficiently.
@@
Figure 1: The Quantum Non-local Neural Network (QNL-Net) comprises a four-qubit circuit composed of three parts: (i) Encoder: To encode classical data into quantum states. (ii) Variational Quantum Circuit (VQC): classically trainable quantum circuit. (iii) Measurement: the circuit is measured at qubit 00 in the Pauli-$Z$ basis. The encoder and the VQC ansatz have $r$ and $D$ repetitions respectively. The Encoder has $4r$ trainable parameters and the VQC has $5D$ trainable parameters (for $n=4$ input qubits).
To encode classical data $X=[y_{0},y_{1},\cdots,y_{n-1}]\in\mathbb{R}^{n}$ into the quantum space, we first write |$\ket{X}$ as the quantum version of $X$ as |$\ket{X}=\ket{y_{0},y_{1},\cdots,y_{n-1}}$. Then, we encode this data as
|$\ket{\psi_{\Phi}}=(\bigotimes^{n}_{k=1}{P(\lambda_{k})}H^{\otimes n})^{r}\ket{X},$ (3)
where $H^{\otimes n}$ is a layer of Hadamard gates acting on all $n$ qubits and $P(\lambda)=\begin{bmatrix}1&0\\ 0&e^{i\lambda}\end{bmatrix}$. Thus, the encoder has $nr$ parameters which can be trained for efficient encoding.
@@
Figure 2: The three ansatzes used as the Variational Quantum Circuits (VQC) in our QNL-Net mechanism. $R_{x},R_{y},R_{z}$ rotation gates represent single-qubit rotations along the $x,y,z$ axes, respectively, with trainable parameters. The strategies for performing entanglement using $C_{X}$ gates (CNOT) are: cyclic pattern (Ansatz-0), reverse linear chain (Ansatz-1), and a mixed pattern (Ansatz-2).
@@
We employ Variational Quantum Circuits (VQC) in our QNL-Net framework, which are classically trainable to optimize parameters using a predefined cost function (Benedetti et al. 2019). This hybrid quantum-classical strategy harnesses the expressive capabilities of quantum circuits while utilizing classical optimization methods for efficient parameter adjustment. Our implementation includes three distinct ansatzes, each incorporating five parameterized rotation gates to provide precise control over the quantum state. These ansatzes feature three $C_{X}$ gates with unique configurations, establishing entanglement-induced correlations among the qubits to ensure that the quantum state captures non-local dependencies within the data. The quantum circuits for these ansatzes are illustrated in Fig. 2. Initially, rotation gates are applied to each qubit in all ansatzes to derive the subsequent quantum state
|$\ket{\psi_{1}}=R_{z}(x_{0})_{q_{0}}R_{y}(\theta_{0})_{q_{1}}R_{y}(\phi_{0})_{q_{2}}R_{x}(g_{0})_{q_{3}}\ket{\psi_{\Phi}},$ (4)
[…]
$R_{y}(\lambda)=\begin{bmatrix}cos(\frac{\lambda}{2})&-sin(\frac{\lambda}{2})\\ sin(\frac{\lambda}{2})&cos(\frac{\lambda}{2})\end{bmatrix},R_{z}(\lambda)=\begin{bmatrix}e^{-i\frac{\lambda}{2}}&0\\ 0&e^{i\frac{\lambda}{2}}\end{bmatrix}.$ (6)
@@
The variation in the ansatzes stems from distinct entanglement strategies outlined below. These configurations ensure comprehensive entanglement among all qubits to capture non-local dependencies across the circuit. In Ansatz-0, the $C_{X}$ gates form a cyclic pattern, establishing a loop of entanglement among the qubits. Ansatz-1 features $C_{X}$ gates forming a reverse linear chain, creating a backward sequential entanglement structure. Ansatz-2 introduces a non-linear and unique mixed pattern entanglement strategy utilizing the $C_{X}$ gates. The resulting quantum states for each respective ansatz are as follows:
|$\ket{\psi_{2}^{[0]}}=C_{X}(q_{1}q_{2})C_{X}(q_{2}q_{3})C_{X}(q_{3}q_{0})\ket{\psi_{1}},$ (7)
|$\ket{\psi_{2}^{[1]}}=C_{X}(q_{3}q_{2})C_{X}(q_{2}q_{1})C_{X}(q_{1}q_{0})\ket{\psi_{1}},$ (8)
|$\ket{\psi_{2}^{[2]}}=C_{X}(q_{1}q_{3})C_{X}(q_{3}q_{2})C_{X}(q_{2}q_{0})\ket{\psi_{1}},$ (9)
where $\text{where }C_{X}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{bmatrix}.$ (10)
@@
This sequence of gates and entanglements in the ansatzes constitutes one layer of the respective ansatz and can also be represented by the unitary operator as
@@
Now, each layer consists of 5 parameters (can be generalized to $n+1$ parameters for $n$ qubits) and can be applied with $D$ repetitions to enhance the expressiveness of the model, resulting in a total of $5D$ trainable parameters from the VQC, and therefore, we obtain the final state of the circuit,
|$\ket{\psi_{s}}=(U_{a}(\theta))^{D}\ket{\psi_{\Phi}},$ (13)
where $a$ is the desired ansatz $a$ = $0,1$, or $2$.
According to the Born rule, measuring any quantum state in the Pauli-$Z$ basis ($\sigma_{z}$) either collapses into the state |$\ket{0}$ with probability $|\alpha|^{2}$ or into the state |$\ket{1}$ with probability $|\beta|^{2}$ (Nielsen and Chuang 2001). In general, the expectation of any observable $\hat{O}$ for a state |$\ket{\psi}$ can be denoted as
@@
$O=\sum_{k}\alpha_{k}P_{k},\hskip 8.53581pt\alpha_{k}\in\mathbb{R},$ (15)
@@
where $\alpha_{k}\in\mathbb{R}$ are real coefficients, and $P_{k}=P_{k_{1}}\otimes P_{k_{2}}\otimes\cdots\otimes P_{k_{n}}$ are tensor products of Pauli operators $P_{k_{i}}\in\{I,X,Y,Z\}$ acting on each of the $n$ qubits. In our mechanism, we measure the circuit in the Pauli-$Z$ computational basis, specifically at one qubit, $q_{0}$, while the rest of the qubits are measured using the Identity ($I$) operation. So, we can denote the measurement of $q_{0}$ for a quantum state of |$\ket{\psi_{s}}$ as $\braket{Z}$ such that,
$\braket{Z}=\bra{\psi_{s}}U^{\dagger}(\theta)Z_{q0}U(\theta)\ket{\psi_{s}},$ (16)
where $\theta$ represents the parameters of the quantum ansatz, $Z_{q_{0}}$ represents the Pauli-$Z$ operator acting on qubit $q_{0}$, $U(\theta)$ is the unitary operator parameterized by $\theta$, and $U^{\dagger}(\theta)$ is the Hermitian adjoint of $U(\theta)$.